OpenStudy (anonymous):

Find the sum of the first 32 terms of the following sequence: -10, -2, 6 . . .

OpenStudy (anonymous):

\[Sn = n/2[2a _{1}+(n-1)d]\]

OpenStudy (anonymous):

n = 32 d = 8 \[a _{1}= -10\]

OpenStudy (anonymous):

\[=32/2[2(-10) + (32-1)8]\]

OpenStudy (anonymous):

= 16[-20 + (31)8]

OpenStudy (anonymous):

= 16 (-20 + 248) = 16 (228) = 3648